OpenStudy (anonymous):

Rewrite 7(40 + 2) as the sum of two products

OpenStudy (mathstudent55):

Use the distributive property: \(a(b + c) = ab + ac\) Follow the pattern above for: \(7(40 + 2)\)

OpenStudy (anonymous):

280+14

OpenStudy (anonymous):

740 + 72?

OpenStudy (mathstudent55):

Both answers above are not the sum of two products. All you need to do is follow the pattern I wrote above and leave it as the sum of products without multiplying it out.

OpenStudy (mathstudent55):

\(7(40 + 2)\) \(= 7 \times 40 + 7 \times 2\)

OpenStudy (anonymous):

@mathstudent55 genius CAP